The perimeter of a rectangle is 220 yards. The length is 56 yards greater than the width. Find the dimensions.
Answer by ikleyn(52946)

.
The perimeter of a rectangle is 220 yards. The length is 56 yards greater than the width.
Find the dimensions.
~~~~~~~~~~~~~~~~~~~~~~~~

    w + L = 220/2

    w + (w+56) = 110

    2w = 110 - 56

    2w = 54

     w = 54/2 = 27.


ANSWER.  The width is 27 yards.  The length is 27+56 = 83 yards.

Solved.